Yen is twice as old as Ben. Four time's Yen's age four years ago is one year less than three times Ben's age five years from now. How old is each now ?

Answer by lwsshak3(11628)

Yen is twice as old as Ben. Four time's Yen's age four years ago is one year less than three times Ben's age five years from now. How old is each now ?
***
let x=Ben's age
2x=Yen's age
..
4 yrs ago:
2x-4=Yen's age
..
5 yrs from now:
x+5=Ben's age
..
4(2x-4)=3(x+5)-1
8x-16=3x+15-1
5x=30
x=6
2x=12
Ben's age=6
Yen's age=12